OpenStudy (idealist10):

Find the volume of the solid under the surface z=xy and above the triangular region in the xy-plane bounded by the lines y=2x, y=-x+6, and y=0.

OpenStudy (idealist10):

@wio @radar @robtobey @Directrix

OpenStudy (anonymous):

draw it

OpenStudy (idealist10):

2x=-x+6 x=2 they intersect at x=2

OpenStudy (anonymous):

Okay, I think that we can start with :\[ 0\leq z \leq xy \]

OpenStudy (anonymous):

I think \(y=0\) means it is \(y\) simple.

OpenStudy (idealist10):

So what are the limits of integration for x and y?

OpenStudy (anonymous):

First solve for \(x\) in terms of \(y\).

OpenStudy (anonymous):

We have \[ x=\frac y2\\x=6-y \]We know at \(0\), that the second equation will have a greater \(x\), so we say: \[ \frac y2 \leq x \leq 6-y \]

OpenStudy (anonymous):

Finally, we know \(y\) starts at \(0\), and solving for the intersection, we get \(y=4\).\[ 0\leq y \leq 4 \]

OpenStudy (idealist10):

Thank you!